### v1

@@ Abstract
Let the measurement settings be chosen from a set of discrete rotations $x\in\{\theta_{A,1},\theta_{A,2},\ldots\theta_{A,n}\}$ and $y\in\{\theta_{B,1},\theta_{B,2},\ldots\theta_{B,n}\}$. Suppose the detector settings are the same, $x=y$, that is to say, the same angle is chosen. We can rotate both angles jointly until we see a perfect anti-correlation at each output. When the photon at $A$ is detected at $a=1$ channel the photon at $B$ is detected at $b=-1$ channel, and vice versa. However from trail to trial the local measurement outcomes are a random binary numbers ±$\pm 1$.
@@ 4 Ringworld: a toy model
Figure 11: An example of the prediction made by a learning machine inside the agent using a long data set of the kind show in Fig (10). Using only four randomly chosen head directions. The blue curve is the ground truth function that agent is trying to learn.
@@
If it receives a random light pulse, the correlation inherent in the ordered pairs $(a_{j},n_{j})$ is contaminated by independent errors in the $n_{j}$. We can represent this by adding/subtracting a small random integer,$e_{j}$ , to each count component
@@
Figure 13: The interior of the disk is composed of a spatially varying refractive index that decreases away from the centre. Red: $\theta=\pi/5$, Blue: $\theta=\pi/4$, Red=$\theta=\pi/3$

### v2

@@ Abstract
Let the measurement settings be chosen from a set of discrete rotations $x\in\{\theta_{A,1},\theta_{A,2},\ldots\theta_{A,n}\}$ and $y\in\{\theta_{B,1},\theta_{B,2},\ldots\theta_{B,n}\}$. Suppose the detector settings are the same, $x=y$, that is to say, the same angle is chosen. We can rotate both angles jointly until we see a perfect anti-correlation at each output. When the photon at $A$ is detected at $a=1$ channel the photon at $B$ is detected at b=1b=-1 channel, and vice versa. However from trail to trial the local measurement outcomes are a random binary numbers $\pm 1$.
@@ 4 Ringworld: a toy model
Figure 11: An example of the prediction made by a learning machine inside the agent using a long data set of the kind show in Fig (10). Using only four randomly chosen head directions. The blue dots represnet the ground truth function that agent is trying to learn.
@@
The case of head angles that are irrational multiples of $\pi$ requires special consideration. Orbits do not close, but almost-returns are certain. The time between them is on average $1/\epsilon$ steps for phase space arc length $\epsilon$ … ‘width’ of agent. For most irrational $\theta/\pi$, the sequence of recurrence times to epsilonepsilon-arcs behaves as:
$Pr(T>t)\sim e^{-\epsilon t}$
where $t$ is the number of bounces.
@@
[…]
@@
[…]